# An algebraic-geometric construction of “lump” solutions of the KP1 equation

### v1

@@ 2. Nodal curves and KP solitons
We begin by setting up some suitable notation for understanding irreducible rational nodal curves and the corresponding soliton KP solutions. We follow notation from [3, 7]. We start from $\mathbb{P}^{1}$ and identify $g$ pairs of points to produce an irreducible rational nodal curve of arithmetic genus $g$. In one of the standard coordinate charts of $\mathbb{P}^{1}$, say the points identified to give the $i$th node are $z=b_{i},c_{i}$. Then a basis for the vector space of dualizing, or Rosenlicht, differentials on the nodal curve is given by
@@
It is well-known (see, for instance, [10, 3, 7, 8]) that the Riemann theta functions on the Jacobians of a family of smooth curves degenerating to one of these rational nodal curves have a limit of the form
@@
(the vector $h$ is analogous to the vector of “Riemann constants” which plays the same role for the theta function from a smooth genus $g$ curve).
@@
In the article [13], Zhang, Yang, Li, Guo, and Stepanyants give examples of degenerations of the $b_{i}$ and $c_{i}$ depending on a parameter $\varepsilon\to 0$ such that the lowest nonzero terms in the series expansion of the $\tau$-function in powers of $\varepsilon$ give rational functions of $x,y,t$ that yield “lump” solutions. However, they do not make the connection with the cuspidal rational curves that are the limits of the nodal rational curves. Hence the connection between their work and ours is that we will use similar degenerations, but we will show explicitly how the limit curve gives a cuspidal rational curve and how the limit theta function corresponds to the polynomial analog of the theta function on the generalized Jacobian of the cuspidal curve.
But in fact, to do this, it will be most convenient to look at the corresponding family of points in the Sato Grassmannian rather than looking explicitly at the limit of the theta function. (That can also be done, of course, but the computations are more awkward.) We will address these points in the next sections.
3. Degenerating to the bi-cuspidal curve, and the polynomial analog of the theta function
@@
$\psi_{4}=\lim_{\varepsilon\to 0}\eta_{4}$ $=\frac{12u^{2}}{(u+1)^{4}}\ du.$
In terms of the original affine coordinate $z=\frac{1}{u}$, these can be written as
$\psi_{1}=\frac{4}{(z-1)^{2}}\ dz,\psi_{2}=\frac{12}{(z-1)^{4}}\ dz,\psi_{3}=\frac{4}{(z+1)^{2}}\ dz,\psi_{4}=\frac{12}{(z+1)^{4}}\ dz,$
@@
By the constructions from [2], the $\langle 2,5\rangle$-cusps correspond to the triangular Young diagrams from the partition $3=2+1$. Moreover, the total degree of the polynomial analog of the theta function is $3+3=6$ in this case. We can find an implicit equation of the $W_{3}$ subvariety of the generalized Jacobian of the cuspidal curve starting from the usual abelian integral parametrization. We take the base point of the abelian integrals as the point $z=0$ or $u=\infty$ on $\mathbb{P}^{1}$, the normalization of the cuspidal curve:
(10) $Z_{j}=\int_{\infty}^{t_{1}}\psi_{j}+\int_{\infty}^{t_{2}}\psi_{j}+\int_{\infty}^{t_{3}}\psi_{j},\quad j=1,\ldots,4.$
Eliminating the $t_{i}$ via a Gröbner basis calculation yields an implicit equation involving a polynomial analog of the theta function:
$-32256Z_{3}-5376Z_{4}=0.$
As expected, the highest-degree term is the $Z_{1}^{3}Z_{3}^{3}$ of total degree $6$.
4. The limiting Sato Grassmannian point and the $\tau$-function
From Theorem 4.11 of [2], to produce a KP2 $\tau$-function from this theta function (up to an exponential factor that does not contribute anything when we apply (2) to produce the actual KP1 solution $u(x,y,t)$), we need to compute a frame for the point of the Sato Grassmannian corresponding to the cuspidal curve. In fact, we have already done the relevant computations needed here, since the crucial part of this comes from the series expansions of the dualizing differentials from (3) above. We have
$\psi_{1}$ $=(4+8u+12u^{2}+16u^{3}+\cdots\ )\ du$
(12) $\psi_{2}$ $=(-12u^{2}-48u^{3}+\cdots\ )\ du$
$\psi_{3}$ $=(4-8u+12u^{2}-16u^{3}+\cdots\ )\ du$
$\psi_{4}$ $=(-12u^{2}+48u^{3}+\cdots\ )\ du$
Hence, taking
$Z_{1}$ $=4x+8iy+12t+\phi_{1}$
(13) $Z_{2}$ $=-12t+\phi_{2}$
$Z_{3}$ $=4x-8iy+12t+\phi_{3}$
$Z_{4}$ $=-12t+\phi_{4}$
and substituting into (11), we obtain what is essentially a $\tau$-function for KP1 solution (note that the complex rescaling of $y$ to $i\cdot y$ has been incorporated here). To clean up the form a bit, we can also divide by the constant factor $4096$. The $\phi_{j}$ are arbitrary constant parameters (analogous to “phase factors” for soliton solutions). They must be chosen appropriately to obtain a regular real solution. In fact, for “most” values of the $\phi_{j}$ the resulting KP1 solutions produced by this recipe will still be non-regular, and they will also take non-real values at some real $(x,y,t)$. We will see how to overcome these difficulties in the next section.
5. Finding a real regular solution
We begin with a simple observation related to the form of (11) and (4). As long as $x,t$ are taken to be real, the only possible terms contributing non-real values in the substituted theta function are those containing odd powers of $y$–that is, $y,y^{3}$, and $y^{5}$–together with arbitrary powers of $x,t$ giving a total degree at most $6$. The coefficient of $y^{5}$ yields the terms
$(-24\phi_{1}+24\phi_{3}+384)i$
$(44-2\phi_{4}+2\phi_{2})i.$
$+(1125/2)t+(249/4)x^{4}+(19521/4)t^{4}.$
Theorem 1.
Proof.
That the $u(x,y,t)$ produced by (2) takes only real values for real $(x,y,t)$ is a consequence of the determination of the $\phi_{j}$ described above and is also clearly visible from the form of (14). To show that this is a regular solution (a “multi-lump”) we will show that the polynomial in (14) is a sum of squares of real polynomials with a nonzero constant term. This will complete the proof. To begin, we note that the polynomial analog of the theta function from (11) above can actually be rewritten in the form
$+36(Z_{1}Z_{3}+10Z_{1}-6Z_{3}-56)(Z_{1}Z_{3}+6Z_{1}-10Z_{3}-56).$
Part of this reflects the general patterns determined in the proof of Theorem 5.3 of [2]. The two factors of degree 3 on the first line are actually the implicit equations of the theta-divisors for the two partial normalizations of the bicuspidal curve, where one of the cusps is smoothed and the other one remains untouched. The other term is of lower total degree 4 and it happens to factor in this way in this case. When the values from (4) with $\phi_{j}$ as above are substituted into this polynomial, the two factors of total degree 3 in the term on the first line become complex conjugates, and the product has the form $(A+iB)(A-iB)=A^{2}+B^{2}$, where $A,B$ are polynomials with real coefficients. The factors on the second line yield
$36((4x+12t+10)^{2}+64y^{2}+4)((4x+12t+6)^{2}+64y^{2}+4)$
Figure 1. The solution (14) at $t=-1.375$.
Figure 2. The solution (14) at $t=-0.9375$.
Figure 3. The solution (14) at $t=-0.4375$.
@@ 6. Comments and Generalizations
On a more hopeful note, much of the construction we have presented, starting from the choice of the pairs of points from (3) for the family of nodal curves generalizes immediately to give families degenerating to cuspidal curves with two $A_{2N}$ double points for all $N\geq 1$. The forms of the dualizing differentials on the cuspidal limits will be parallel, and the results of [2] yield a similar factorization of the leading terms of the polynomial analog of the theta function on the cuspidal curve. Theorem 4.11 of [2] applies in general as here to give KP1 solutions.
@@
In additional support of this conjecture, we include a plot of a similar solution constructed from a bicuspidal curve with two $\langle 2,7\rangle$ cusps ($A_{6}$ double points) following the same plan as that used in the calculations reported above. This comes from a family of rational nodal curves as in (3), but with
@@
The polynomial analog of the theta function in this case has degree 12. The polynomials involved are too complicated to be readily understandable, so they are omitted. The pattern established in Theorem 5.3 of [2] is clear, though. The highest degree term in the polynomial analog of the theta function is $Z_{1}^{6}Z_{4}^{6}$ since the corresponding partition for each cusp is the triangular $6=3+2+1$.
@@
The obstacle that we have not overcome as of yet is proving that there will always be choices of the “phase factors” $\phi_{j}$ as above that produce real regular KP1 solutions. This will require more detailed information about the form of the terms of lower total degree in the polynomial analog of the theta function on the cuspidal limit curve. This seems somewhat similar to, but more complicated than, the form seen in (15) when $N\geq 3$. We hope to return to this in the future.

### v2

@@ 6. Nodal curves and KP solitons
To show how our method relates with other previous work producing multi-lump solutions, we will now describe an alternative way to derive the KP1 solution corresponding to (11). We begin by setting up some suitable notation for understanding irreducible rational nodal curves and the corresponding soliton KP solutions. We follow notation from [3, 7]. We start from $\mathbb{P}^{1}$ and identify $g$ pairs of points to produce an irreducible rational nodal curve of arithmetic genus $g$. In one of the standard coordinate charts of $\mathbb{P}^{1}$, say the points identified to give the $i$th node are $z=b_{i},c_{i}$. Then a basis for the vector space of dualizing, or Rosenlicht, differentials on the nodal curve is given by
@@
It is well-known (see, for instance, [10, 3, 7, 8]) that the Riemann theta-functions on the Jacobians of a family of smooth curves degenerating to one of these rational nodal curves have a limit of the form
@@
(the vector $h$ is analogous to the vector of “Riemann constants” which plays the same role for the theta-function from a smooth genus $g$ curve).
@@
In the article [14], Zhang, Yang, Li, Guo, and Stepanyants give examples of degenerations of the $b_{i}$ and $c_{i}$ depending on a parameter $\varepsilon\to 0$ such that the lowest nonzero terms in the series expansion of the $\tau$-function in powers of $\varepsilon$ give rational functions of $x,y,t$ that yield “lump” solutions. However, they do not make the connection with the cuspidal rational curves that are the limits of the nodal rational curves. Hence the connection between their work and ours is that we will use similar degenerations, but we will show explicitly how the limit curve gives a cuspidal rational curve and how the limit theta-function corresponds to the polynomial analog of the theta-function on the generalized Jacobian of the cuspidal curve.
But in fact, to do this, it will be most convenient to look at the corresponding family of points in the Sato Grassmannian rather than looking explicitly at the limit of the theta-function. (That can also be done, of course, but the computations are more awkward.) We will address these points in the next sections.
7. Degenerating to the bi-cuspidal curve, and the polynomial analog of the theta-function
@@
$\psi_{4}=\lim_{\varepsilon\to 0}\eta_{4}$ $=\frac{-12u^{2}}{(u+1)^{4}}\ du.$
These are the same as the differentials from (4). In terms of the original affine coordinate $z=\frac{1}{u}$, these can be written as
$\psi_{1}=\frac{-4}{(z-1)^{2}}\ dz,\psi_{2}=\frac{12}{(z-1)^{4}}\ dz,\psi_{3}=\frac{-4}{(z+1)^{2}}\ dz,\psi_{4}=\frac{12}{(z+1)^{4}}\ dz,$
@@
Then the rest of the derivation follows as before, in particular the use of Theorem 3 to derive the KP tau-function.
@@ 8. Comments and Generalizations
On a more hopeful note, much of the construction we have presented, starting from the choice of the pairs of points from (7) for the family of nodal curves generalizes immediately to give families degenerating to cuspidal curves with two $A_{2N}$ double points for all $N\geq 1$. The forms of the dualizing differentials on the cuspidal limits will be parallel, and the results of [2] yield a similar factorization of the leading terms of the polynomial analog of the theta-function on the cuspidal curve. Theorem 4.11 of [2] applies in general as here to give KP1 solutions.
@@
In additional support of this conjecture, we include a plot of a similar solution constructed from a bicuspidal curve with two $\langle 2,7\rangle$ cusps ($A_{6}$ double points) following the same plan as that used in the calculations reported above. This can be obtained from a family of rational nodal curves as in (7), but with
@@
The polynomial analog of the theta-function in this case has degree 12. The polynomials involved are too complicated to be readily understandable, so they are omitted. The pattern established in Theorem 5.3 of [2] is clear, though. The highest degree term in the polynomial analog of the theta-function is $Z_{1}^{6}Z_{4}^{6}$ since the corresponding partition for each cusp is the triangular $\lambda=(3,2,1)$.
@@
Finally, in Figure 5, we display a time snapshot of another five-lump solution constructed from a rational curve with two cusps having semigroup $\langle 3,4\rangle$.
Figure 5. A KP1 solution with 5 lumps.
The corresponding partitions are both $\lambda=(3,1,1)$ and the polynomial analog of the theta-function has degree $10=2\cdot(3+1+1)$.
The obstacle that we have not overcome as of yet to generalize these results is proving that there will always be choices of the “phase factors” $\phi_{j}$ as above that produce real regular KP1 solutions. This might be obtained with more detailed information about the form of the terms of lower total degree in the polynomial analog of the theta-function on the bicuspidal curve. This seems somewhat similar to, but more complicated than, the form seen in (12) when the degree is larger. We hope to return to this in the future.